Packaging Unit Speed and Damage Probability

XAT 2010 · Quantitative Ability & Data Interpretation · Hard · Quantitative Ability

The supervisor of a packaging unit of a milk plant is being pressurised to finish the job closer to the distribution time, thus giving the production staff more leeway to cater to last minute demand. He has the option of running the unit at normal speed or at 110% of normal - "fast speed". He estimates that he will be able to run at the higher speed for 60% of the time. The packet is twice as likely to be damaged at the higher speed which would mean temporarily stopping the process. If a packet on a randomly selected packaging runs has probability of 0.112 of damage, what is the probability that the packet will not be damaged at normal speed?

  1. A.

    0.81

  2. B.

    0.93

  3. C.

    0.75

  4. D.

    0.60

  5. E.

    None of the above

Answer

B

Explanation

Let mm be rate at normal speed. Output in time tt at normal speed = 0.40×m⋅t=0.40mt0.40 \times m \cdot t = 0.40 mt. At fast speed (110%): 0.60×1.10m⋅t=0.66mt0.60 \times 1.10 m \cdot t = 0.66 mt. Proportion of output at normal speed = 0.34mtmt=0.34\frac{0.34 mt}{mt} = 0.34. Proportion at fast speed = 0.660.66. Let pp be probability of damage at normal speed. At fast speed, damage probability = 2p2p. Total damage probability = 0.34p+0.66(2p)=1.66p=0.112  ⟹  p=0.1121.66=0.0670.34p + 0.66(2p) = 1.66p = 0.112 \implies p = \frac{0.112}{1.66} = 0.067. Probability of NOT damaged at normal speed = 1−0.067=0.933≈0.931 - 0.067 = 0.933 \approx 0.93.

Related Quantitative Ability questions

Practise this under exam conditions

Sign in to solve it with a live timer, the on-screen CAT calculator, and streak and accuracy tracking across every question you attempt.

Solve in the workspace